How much energy does it take to make 500 g of solid lead at 21°C and turn it entirely to liquid lead?
1
Expert's answer
2020-05-29T10:07:34-0400

The total amount of energy required to turn solid lead entirely to liquid lead consists of two parts:


"Q_{tot} = Q_{heat} + Q_{melt}"

where "Q_{heat}" is the energy required to heat lead up to the melting point and "Q_{melt}" is the energy requred to melt the given amount of lead.


The first term:


"Q_{heat} = mc(T_{melt}-T_0)"

where "m = 0.5 kg" is the mass of lead, "c = 127.5 \dfrac{J}{kg\cdot K}" is the specific heat capacity of lead,"T_{melt} = 600.6 K" is the melting point of lead and "T_0 = 294.2 K" is the initial temperature (21°C).

Substitute numerical values:


"Q_{heat} = mc(T_{melt}-T_0) = 0.5\cdot 127.5\cdot (600.6 - 294.2) = 19.533 kJ"


The second term:


"Q_{melt} = m\lambda"

where "\lambda = 25 kJ/kg" is the specific heat of fusion of lead.

Substitute numerical values:


"Q_{melt} = m\lambda = 0.5\cdot 25 = 12.5 kJ"

The total energy, thus, will be:


"Q_{tot} = 19.533 kJ + 12.5 kJ = 32.033 kJ \approx32 kJ"

Answer. It takes 32 kJ to make 500 g of solid lead at 21°C and turn it entirely to liquid lead.